OpenStudy (help!!!!):

The Perez family has 100 more shares of stock b than of stock A. The current price per share of stock A is 18.50 and of stock B is $24.75. At most how many shares of each do they have if the average price per share is greater than $22?

OpenStudy (help!!!!):

@ganeshie8 @Loser66

OpenStudy (loser66):

Any idea?

OpenStudy (help!!!!):



OpenStudy (help!!!!):

This is what I got so far

OpenStudy (loser66):



OpenStudy (loser66):

very good, go ahead, solve for x

OpenStudy (help!!!!):

x is greater than 366.6

OpenStudy (help!!!!):

so would my answers be a=366 and b=466?

OpenStudy (help!!!!):

ops I mean x is less than 366.6

OpenStudy (loser66):

\(\dfrac{18.50 x +24.75 (x+100)}{2x+100}\geq 22\) \(18.50x +24.75 x +2475 \geq 22(2x+100)\geq 44x +2200\) \(43.25x +2475 \geq 44x +2200\) \(0.75 x \leq 475\) \(x \leq 633.33\)

OpenStudy (help!!!!):

isnt -2475+2200= -275?

OpenStudy (loser66):

oh, yeah, my bad!! I am sorry

OpenStudy (help!!!!):

Anyways thanks for the help and clarification!